OpenStudy (cloverracer):

What are the coordinates of the vertex of the graph of the function y = -3x^2 - 12x + 3? a. (-2,29) b. (2, 15) c. (2, -9) d. (2, -15)

OpenStudy (cloverracer):

For the x-coordinate I got, x = 2. But for some reason I got -57 as my y-coordinate?

OpenStudy (anonymous):

ok so were is the graph

OpenStudy (cloverracer):

There is no graph. You have to find the coordinates.

OpenStudy (anonymous):

ok cool

OpenStudy (cloverracer):

@cwrw238 What did I do wrong?

OpenStudy (anonymous):

try negative 2

OpenStudy (cwrw238):

i'm just checking it out

OpenStudy (cwrw238):

yes its -2

OpenStudy (cloverracer):

how?

OpenStudy (cwrw238):

convert it to vertex form and you'll see - the y coordinate is 15

OpenStudy (anonymous):

at vertex slope of graph is 0 put dy/dx=0 you will get x

OpenStudy (cwrw238):

yes - you can do it by differentiation

OpenStudy (anonymous):

and after that put value of x in y = -3x^2 - 12x + 3 you will get corresponding y i get pair (-2,15)

OpenStudy (cwrw238):

maybe clover has 'nt done calculus yet have you familiar with differentiation clover?

OpenStudy (cloverracer):

nope lol im in algebra I and just learning quadratic functions

OpenStudy (cloverracer):

I got my x-coordinate by using the equation of the line of symmetry which is x = -b/2a

OpenStudy (cloverracer):

if I use 2 as my x-coordinate I get -57 as my y-coordinate. if I use -2 as my x-coordinate I get -33?

OpenStudy (cwrw238):

Ok we have -3x^2 - 12x + 3 = -3(x^2 + 4x + 1) = -3[(x +2)^2 -4 + 1) this gives the x coordinate of the vertex = -2 now plug -2 into the original equation and we get y = 5 vertex is at (-2,15)

OpenStudy (cloverracer):

oooohhh okay! Thanks for the help! ((:

OpenStudy (cwrw238):

thats ok - its not one of the options but it is correct

OpenStudy (cwrw238):

-b/2a gives you -2 - (-12) / -6 = 12/-6 = -2